# The $\beta$-decay spectrum of Tritiated graphene: combining nuclear quantum mechanics with Density Functional Theory

### v1

@@ III.3 The Tritium parallel potential
Importantly, in the region of about $1.5$ around the equilibrium point, the potential is very similar in all three directions, due to the high degree of isotropy induced by the hexagonal symmetry of the graphene lattice. In this region, the curves corresponding to $z<z_{0}$ (black line) display repulsion at $x=y=0$, since the C-T bond is compressed. Conversely, all the other curves display an attractive minimum, which becomes shallower as the distance between Tritium and graphene increases, as expected.
For displacements larger than $1.5$, instead, large anisotropies appear. In particular, $U_{x}$ presents a large potential barrier (see, again, Fig. 3, panel (d)), since moving in that direction the Tritium encounters another one bound in meta position at $x=2.5$. In the case of $U_{y}$ and $U_{30}$, instead, the nearest Tritium is located at $4.3$, and therefore the barrier is considerably farther away, out of the range probed here. These results, together with the potential $U_{z}$ described in the previous section, give a good description of the complete potential, $U_{\rm T}(\bm{x})$, for $100\%$ loading. This will serve as the basis to compute the initial state wave function to be used in the $\beta$-decay rate, as anticipated in Section II and detailed in Section V.
@@ IV Helium right after the decay
We now move to the analysis of the final state of the reaction in Eq. (1): Helium in interaction with the rest of the substrate immediately after the decay. Near the end-point, the $\beta$-decay happens over a time scale of about $t_{\beta}\sim 10^{-18}$ s, which is conservatively estimated as the time it takes for the $\beta$-electron, with momentum $p_{\beta}\sim 100$ keV, to escape the Helium atom, whose radius is $r_{\rm He}\sim 1$.
Immediately after the decay, the electronic structure will still be the one corresponding to the ground state of the system before the decay. However, for the new system, which now has a Helium nucleus, this will not correspond to the ground state anymore, but it will rather be a combination of excited electronic states (named He in panel (c) of Fig. 1). The time it takes for this state to relax down to the ground state is very variable, depending on the specific mechanisms of non-adiabatic relaxation, but it is generally not smaller than $t_{\rm r}\sim 10^{-16}-10^{-14}$ s Marciniak et al. (2019); Tichauer et al. (2022). We can then safely assume $t_{\beta}\ll t_{\rm r}$. This circumstance evidently prevents the use of the BO approximation, which requires that the electronic structure corresponds to an eigenstate of the current nuclear configuration.
In this context, one often uses the so-called sudden approximation (e.g., Sakurai and Napolitano, 2020; Saenz and Froelich, 1997). In our case, such an approximation is valid because, on the time scale relevant for the decay, the electronic structure of the system does not have enough time to relax, thus remaining the same as it was right before the transition. However, the implementation of this approximation in our context requires to abandon the standard DFT scheme. Specifically, referring to Eq. (7), the electronic density, $n(\bm{x})$, should not be calculated self-consistently given the new nuclear configuration (with a Helium), but must be kept the same as it was before the decay. As consequence, the term $E^{\rm int}[n]$ in Eq. (7) is the same as the one evaluated before decay. Conversely, the terms $E^{\rm NN}(\{\bm{X}\})$ and $E^{e{\rm N}}[n,\{\bm{X}\}]$ do vary, since the nuclear charge has now changed. The calculation of these non-self-consistent terms are not implemented in the standard DFT, and therefore require specific post-processing tasks, which we describe in the SI, Section  SI.I.E.
Even within the sudden approximation, the final Helium potential depends on the choice we make for the electronic density, $n$ [^3] As before, the potentials also depend on how other nuclei are treated. For this case, we use the prescription where all the other are kept forzen. In particular, we implement two possible choices, whose results are reported in Fig. 4, panel (a).
@@
In Fig. 4 we also report the Helium potential obtained within the full adiabatic approximation (i.e., the standard BO framework). In this case, two electrons follow the Helium nucleus, which detaches from the substrate as a neutral atom. As a noble element, an unbound Helium atom is very stable, and therefore its potential is repulsive.
Figure 4: Orthogonal Helium potential in sudden and semi-sudden approximation (respectively, SA and SSA, in red and magenta), and in the adiabatic (BO) approximation (blue, the dashed part is for $z<z_{0}$). The system is considered electrically isolated during decay, and therefore the total charge is +1. Calculations are performed keeping all nuclei frozen, except the one under consideration. The corresponding parallel potentials are reported in Fig.  SI.3. The irregularities in the profiles of the SSA and BO curves, at about 2.7 Å, are the residuals of the barrier inherited by the Tritium potential and the van der Waals attraction well, respectively.
@@ V The $\beta$-decay rate
Our starting point to determine the wave functions, both initial and final, is the generic form of the Schrödinger equation for the nuclei participating to the decay process,
@@
Figure 5: Electron $\beta$-decay spectra obtained for each of the three schemes employed to define the final Helium potential. The initial Tritium potential is taken to be the one corresponding to 100% loading, computed in the prescription where all other nuclei are kept fixed (see Section III). The total Tritium mass is taken to be $M_{\rm tot}=1\text{ }\upmu\text{g}$. Upper panels: Total rate for $m_{\nu}=0.2$ eV. We also show the partial contribution for some selected final states (or groups of final states), as indicated directly in the figures. Lower panels: Total rates in a region much closer to the end-point, as computed for different values of the neutrino mass.
We first recall that in the bound region, i.e. for energies $\varepsilon_{A}<U_{f}$, the potentials reported in Fig. 3 (and Fig.  SI.3) look approximately the same if computed as a function of the three parallel directions we probed. Large anisotropies only appear for distances sufficiently away from the minimum, as commented in Section III.3. These regions however, give a minor contribution to the matrix element in Eq. (5), since the initial Tritium wave function is exponentially localized in the vicinity of the minimum. This allows to approximate the Tritium and Helium potentials as isotropic on the graphene plane, $U(\bm{x})\simeq U(r,z)$. If one is considering a bound state, the dependence of its wave function on the polar angle defined on the graphene plane, $\theta$, is then completely determined to be,
$\psi_{n\ell}(\bm{x})=\chi_{n\ell}(r,z)\,\frac{e^{i\ell\theta}}{\sqrt{2\pi}}\,,$ (13)
with $\ell\in\mathds{Z}$. The quantum number $n$ labels the bound state solution to the remaining Schrödinger equation,
$\begin{split}-\frac{1}{2m}&\left[\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial}{\partial r}\right)-\frac{\ell^{2}}{r^{2}}+\frac{\partial^{2}}{\partial z^{2}}\right]\chi_{n\ell}\\ &\qquad\qquad\qquad+U(r,z)\,\chi_{n\ell}=\varepsilon_{n\ell}\,\chi_{n\ell}\,.\end{split}$ (14)
@@
$\psi_{i,\mathrm{T}}(\bm{x})=\frac{\chi_{00}^{\mathrm{T}}(r,z)}{\sqrt{2\pi}}\,,\quad\psi_{f,\mathrm{He}}(\bm{x})=\chi_{n\ell}^{\mathrm{He}}(r,z)\,\frac{e^{i\ell\theta}}{\sqrt{2\pi}}\,.$ (15)
For final states where the Helium is, instead, in the continuous spectrum (i.e., when it is freed from the graphene right after the decay), we simply approximate the wave function by a plane wave. The details on how we solve the Schrödinger equation, and how we compute the event rates, are reported extensively in the SI, Section  SI.II.
@@
However, the rates obtained with the different schemes differ in two ways. First of all, the end-point energy for the sudden scheme is substantially larger than that in the semi-sudden scheme. The maximum electron energy, in fact, is given by,
$\begin{split}K_{\beta}^{\mathrm{end}}\equiv{}&Q-m_{\nu}-\big{(}\big{|}\varepsilon_{00}^{\mathrm{T}}\big{|}-\big{|}\varepsilon_{00}^{\mathrm{He}}\big{|}\big{)}\\ ={}&Q-m_{\nu}-\big{(}\big{|}E_{b}^{\rm T}\big{|}-\big{|}E_{b}^{\rm He}\big{|}\big{)}\\ &+\frac{1}{2}\big{(}\omega_{\rm t}^{\rm T}-\omega_{\rm t}^{\rm He}\big{)}+\big{(}\omega_{\rm p}^{\rm T}-\omega_{\rm p}^{\rm He}\big{)}\,,\end{split}$ (16)
where $Q\equiv m_{\rm T}-m_{\rm He}-m_{e}$ is the $Q$-value determined by the Tritium and Helium nuclear masses, and $\varepsilon_{00}^{\rm T}$ and $\varepsilon_{00}^{\rm He}$ are the ground state binding energies of Tritium and Helium, respectively. In the last line we have expressed these binding energies in terms of the parameters of the potentials, in particular of the depth of the potentials, $E_{b}$, the fundamental frequency in the direction transverse to the graphene, $\omega_{\rm t}$, and the fundamental frequency in the direction parallel to the graphene, $\omega_{\rm p}$. These are all quantities that can be determined for different configurations. Now, the Helium potential in the sudden case is much deeper than that in the semi-sudden one, resulting in a binding energy that is about 42 eV larger (in magnitude), thus inducing the shift in the end-point energy.
This is also the reason for the second qualitative difference: the separation between the end-point and the onset of the contribution coming from the Helium continuous final states is much widely separated in the sudden scheme than in the semi-sudden one. In fact, the energy at which the Helium can be freed from the graphene layer is determined solely by the depth of the initial Tritium potential, as it is given by,
$\begin{split}K_{\beta}^{\rm free}\equiv{}&Q-m_{\nu}-\big{|}\varepsilon_{00}^{\rm T}\big{|}\\ ={}&Q-m_{\nu}-\big{|}E_{b}^{\rm T}\big{|}+\frac{1}{2}\omega_{\rm t}^{\rm T}+\omega_{\rm p}^{\rm T}\,.\end{split}$ (17)
Consequently, the larger is $\big{|}\varepsilon_{00}^{\rm He}\big{|}$ the more separated the contribution of the continuum from the end-point is. [^4] Indeed, for the sudden case, the onset of the continuous states is so far from the end-point that it does not show in our plot. For energies sufficiently far away from the end-point (but not so far away that the neutrino momentum can no longer be neglected), the contribution coming from continuous final states might dominate the total decay rate, and consequently the experimental statistics.
@@
Lastly, in the adiabatic scheme the final Helium potential is purely repulsive and, therefore, does not allow for bound states. Consequently, the electron spectrum gets contributions only from the continuous final states. Since we approximate the free Helium wave function by a simple plane wave, the corresponding contribution in the event rate is identical in all three instances we considered: sudden, semi-sudden and adiabatic.
@@ VI The fate of Helium
Fig. 6, panel (a) reports the dynamics of the system for the first picosecond, starting from the equilibrium configuration before the decay with null velocities (i.e., roughly zero temperature). We have studied the dynamics of the system for two different configurations: electronically isolated from the outside, and grounded. The red curves in Fig. 6 correspond to the first instance, in which the system has a total charge of +1, while the black curves correspond to the second instance, where the system is neutralized by gaining an electron from the outside. We find that the Helium detaches in few fs with an energy gain for the system $>2$ eV, as expected (numerical values of these energies for different cases, are reported in the SI, Fig.  SI.1). The oscillations which are visible in Fig. 6 are due to the vibrational excitations induced by the recoil of the Carbon site on which the decay has happened, shown in the structure reported as inset. When the system is neutral (black lines), the behavior is similar, but the energy gain is nearly double, as a consequence of the larger stability of neutral system.
Just after release, the total kinetic energy gets a contribution due to the velocity acquired by the Helium, which is $v_{\rm He}\sim 2.9\times 10^{-5}\simeq 0.088$/fs, for the overall charged case, and $v_{\rm He}\sim 5\times 10^{-5}\simeq 0.15$/fs, when the system is electrically neutral. These values correspond to a Helium kinetic energy of about $1.2$ eV and $3.5$ eV, respectively. As the system evolves, Helium bumps on the graphene layers, and gradually distributes its energy to the substrate over a time scale of ps. This is shown in the upper part of Fig. 6, panel (b), where we also show a prototypical trajectory in the inset. As a consequence, the average energy of the system increases of a few eV.
@@ VII Limitations of our approach and theory systematics
As already explained, to successfully constrain the possible values of the neutrino mass it is of utmost importance to provide a reliable theoretical prediction for the electron’s $\beta$-decay spectrum. The present work is a first step in this direction, as we connect the low energy properties of Tritiated graphene to the shape exhibited by the spectrum. The framework we are currently working in, however, has some limitations, which we now discuss.
@@ VIII Conclusions and Perspectives
The issues raised in the previous section could all be summarized in the consideration that the established and more commonly used theoretical scheme generally rely on some sort of separation of energy or time scale, which in turns allows to decouple a subset of degrees of freedom from others, facilitating the approach to the problem. In this work, conversely, we aimed at addressing a question which is at the connection point of high and low energy scales, overlapping on different time scales, bridging solid state with nuclear physics. Therefore the approaches relying on the separation of degrees of freedom, although can be used as the starting point, needs to be reconsidered and possibly revised.
@@
However, a stronger assumption we made is about the decoupling of the electronic coordinates from the nuclear ones. We used it in two rather opposite declinations: the BO approximation, used before and long after the decay, assuming total adiabaticity of the electrons with respect to the nuclear motion, and the sudden approximation, assuming on the contrary frozen electronic structure. Between these two extremes, a number of theoretical schemes are currently considered to address the relaxation of electronic dynamics during nuclear motion, especially when the latter involves light elements, where quantum effects are more evident. Aside the evolution of electronic excited states (e.g., within the Time Dependent DFT schemes Marques et al. (2006)) one must also consider the coupled nuclear motion, either within a semi-classical scheme Li et al. (2005), or within the path-integral scheme Kim et al. (1997), if a full quantum accuracy is required. While the latter is still too computationally heavy to address the system of interest, an interesting approach has recently emerged for a quantum perturbative anharmonic treatment with a limited computational cost Monacelli and Mauri (2021).
The analysis of the model system also requires some consideration. Here we have considered only a few possible loading levels, and with symmetric distributions. In a real system, the situation is rather different. While the loading can be somehow controlled and measured, the local Tritium distribution at given loading is not. On the other hand, our results (and others’ too Delfino et al. (2024); Rossi et al. (2015); Camiola et al. (2015)) indicate that the parameters of the potentials (binding and barrier energies) are rather dependent on the local level of loading. Therefore, if one wants to achieve a realistic determination of the parameters of the potential (and in turn of the $\beta$-decay spectrum) one should consider the statistical distribution of possible environments of Tritium/Helium on graphene.
@@
All the aspects above will eventually make up the ingredients leading to the determination of the $\beta$-electron spectrum, which should be compared to data to constrain the value of the neutrino mass. To this end, it will be crucial to develop the above proposals, and possibly any other that might be relevant, with a quantifiable, controlled accuracy. This will allow to provide a theoretical systematics, to be fed into the experimental analysis.

### v2

@@ III.3 The Tritium parallel potential
Importantly, in the region of about $1.5$ Å around the equilibrium point, the potential is very similar in all three directions, due to the high degree of isotropy induced by the hexagonal symmetry of the graphene lattice. In this region, the curves corresponding to $z<z_{0}$ (black line) display repulsion at $x=y=0$, since the C-T bond is compressed. Conversely, all the other curves display an attractive minimum, which becomes shallower as the distance between Tritium and graphene increases, as expected.
For displacements larger than $1.5$ Å, instead, large anisotropies appear. In particular, $U_{x}$ presents a large potential barrier (see, again, Fig. 3, panel (d)), since moving in that direction the Tritium encounters another one bound in meta position at $x=2.5$ Å. In the case of $U_{y}$ and $U_{30}$, instead, the nearest Tritium is located at $4.3$ Å, and therefore the barrier is considerably farther away, out of the range probed here. These results, together with the potential $U_{z}$ described in the previous section, give a good description of the complete potential, $U_{\rm T}(\bm{x})$, for $100\%$ loading. This will serve as the basis to compute the initial state wave function to be used in the $\beta$-decay rate, as anticipated in Section II and detailed in Section V.
@@ IV Helium right after the decay
We now move to the analysis of the final state of the reaction in Eq. (1): Helium in interaction with the rest of the substrate immediately after the decay. Near the end-point, the $\beta$-decay happens over a time scale of about $t_{\beta}\sim 10^{-18}$ s, which is conservatively estimated as the time it takes for the $\beta$-electron, with momentum $p_{\beta}\sim 100$ keV, to escape the Helium atom, whose radius is $r_{\rm He}\sim 1$ Å.
Immediately after the decay, the electronic structure will still be the one corresponding to the ground state of the system before the decay. However, for the new system, which now has a Helium nucleus, this will not correspond to the ground state anymore, but it will rather be a combination of excited electronic states (named He∗ in panel (c) of Fig. 1). The time it takes for this state to relax down to the ground state is very variable, depending on the specific mechanisms of non-adiabatic relaxation, but it is generally not smaller than $t_{\rm r}\sim 10^{-16}-10^{-14}$ s Marciniak et al. (2019); Tichauer et al. (2022). We can then safely assume $t_{\beta}\ll t_{\rm r}$. This circumstance evidently prevents the use of the BO approximation, which requires that the electronic structure corresponds to an eigenstate of the current nuclear configuration.
In this context, one often uses the so-called sudden approximation (e.g., Sakurai and Napolitano, 2020; Saenz and Froelich, 1997). In our case, such an approximation is valid because, on the time scale relevant for the decay, the electronic structure of the system does not have enough time to relax, thus remaining the same as it was right before the transition. However, the implementation of this approximation in our context requires to abandon the standard DFT scheme. Specifically, referring to Eq. (7), the electronic density, $n(\bm{x})$, should not be calculated self-consistently given the new nuclear configuration (with a Helium), but must be kept the same as it was before the decay. As a consequence, the term $E^{\rm int}[n]$ in Eq. (7) is the same as the one evaluated before decay. Conversely, the terms $E^{\rm NN}(\{\bm{X}\})$ and $E^{e{\rm N}}[n,\{\bm{X}\}]$ do vary, since the nuclear charge has now changed. The calculation of these non-self-consistent terms are not implemented in the standard DFT, and therefore require specific post-processing tasks, which we describe in the SI, Section  SI.I.E.
Even within the sudden approximation, the final Helium potential depends on the choice we make for the electronic density, nn. [^4] As before, the potentials also depend on how other nuclei are treated. For this case, we use the prescription where all the other are kept forzen. In particular, we implement two possible choices, whose results are reported in Fig. 4, panel (a).
@@
In Fig. 4 we also report the Helium potential obtained within the full adiabatic approximation (i.e., the standard BO framework). In this case, two electrons follow the Helium nucleus, which detaches from the substrate as a neutral atom. As a noble element, an unbound Helium atom is very stable, and therefore its potential is repulsive. The vertical displacement of the adiabatic curve with respect to the sudden/semi-sudden at their minima is due to the electronic structure relaxation, and turns out to be approximately $18$ eV. The figure also includes the adiabatic potential for Tritium. Within DFT calculations, the absolute energy reference of systems with different constituents is generally different. Consequently, to correctly align the Tritium potential to the others, we used as reference the Coulomb tails of potentials for the Helium and Tritium nuclei. These are aligned using the respective sudden potentials (that for Tritium is not shown here, but has been computed for this purpose). More details are given in the SI, section SI.I.F. This sets the energy difference between the minimum of the Tritium potential, and the minimum of the sudden/semi-sudden potentials, to be $U_{0}^{\rm T}-U_{0}^{\rm SA/SSA}\simeq 29$ eV.
Figure 4: Orthogonal Helium potential in sudden and semi-sudden approximation (respectively, SA and SSA, in blue and red), and in the adiabatic (BO) approximation (green). The system is considered electrically isolated during decay, and therefore the total charge is +1, although in the three cases the Helium is extracted in different ionic states, as indicated. The corresponding parallel potentials are reported in Fig.  SI.4. For comparison, the Tritium potential is also reported, aligned to a common energy frame as explained in the text and in the SI, section SI.I.F.
@@ V The $\beta$-decay rate
Our starting point to determine the wave functions, both initial and final, is the generic form of the SchrĂśdinger equation for the nuclei participating in the decay process,
@@
Figure 5: Electron $\beta$-decay spectra obtained for each of the three schemes employed to define the final Helium potential. The initial Tritium potential is taken to be the one corresponding to 100% loading, computed in the prescription where all other nuclei are kept fixed (see Section III). The total Tritium mass is taken to be $M_{\rm tot}=1\text{ }\upmu\text{g}$. The electron’s kinetic energy is measured with respect to the end-point of a Tritium nucleus in vacuum and for a massless neutrino, $K_{0}$. Upper panels: Total rate for $m_{\nu}=0.2$ eV. We also show the partial contribution for some selected final states (or groups of final states), as indicated directly in the figures. Lower panels: Total rates in a region much closer to the end-point, as computed for different values of the neutrino mass.
We first recall that in the bound region, i.e. for energies $\varepsilon_{A}<U_{f}$, the potentials reported in Fig. 3 (and Fig.  SI.4) look approximately the same if computed as a function of the three parallel directions we probed. Large anisotropies only appear for distances sufficiently away from the minimum, as commented in Section III.3. These regions however, give a minor contribution to the matrix element in Eq. (5), since the initial Tritium wave function is exponentially localized in the vicinity of the minimum. This allows to approximate the Tritium and Helium potentials as isotropic on the graphene plane, $U(\bm{x})\simeq U(r,z)$. If one is considering a bound state, the dependence of its wave function on the polar angle defined on the graphene plane, $\theta$, is then completely determined to be,
[…]
@@
[…]
For final states where the Helium is, instead, in the continuous spectrum (i.e., when it is freed from the graphene right after the decay), we simply approximate the wave function by a plane wave. The details on how we solve the SchrÜdinger equation, and how we compute the event rates, are reported extensively in the SI, Section  SI.II.
@@
Nonetheless, the rate obtained with different schemes, differ in a few ways. First of all, when final Helium bound states are allowed, the maximum electron energy is,
[…]
where $Q\equiv m_{\rm T}-m_{\rm He}-m_{e}$ is the $Q$-value determined by the Tritium and Helium nuclear masses, and $\varepsilon_{00}^{\rm T}$ and $\varepsilon_{00}^{\rm He}$ are the energies of the ground state of Tritium and Helium, respectively. In the last line we have expressed these energies in terms of the parameters of the potentials, in particular, the value of their minima, $U_{0}$, the fundamental frequency in the direction transverse to the graphene, $\omega_{\rm t}$, and the fundamental frequency in the direction parallel to the graphene, $\omega_{\rm p}$. Now, while the minimum of the potential is the same for both the sudden and semi-sudden schemes, the fundamental frequencies are not, as they depend on the actual shape of the potential close to its minimum.
The two schemes also differ in the separation between the end-point and the onset of the contribution coming from the Helium continuous final states. In fact, the energy at which the Helium can be freed from the graphene layer is determined by the difference between the energy of the initial bound Tritium, and the value taken by the Helium potential at infinity, $U_{\infty}^{\rm He}$. Specifically, it is given by,
$\begin{split}K_{\beta}^{\rm free}\equiv{}&Q-m_{\nu}+\varepsilon_{00}^{\rm T}-U_{\infty}^{\rm He}\\ ={}&Q-m_{\nu}+U_{0}^{\rm T}+\frac{1}{2}\omega_{\rm t}^{\rm T}+\omega_{\rm p}^{\rm T}-U_{\infty}^{\rm He}\,.\end{split}$ (17)
Since the value of the potential at infinity is much lower for the semi-sudden approximation than it is for sudden one, the onset of the continuum in the former is substantially closer to the end-point than it is for the latter. [^5] Indeed, for the sudden case, the onset of the continuous states is so far from the end-point that it does not show in our plot. For energies sufficiently far away from the end-point (but not so far away that the neutrino momentum can no longer be neglected), the contribution coming from continuous final states might dominate the total decay rate, and consequently the experimental statistics.
@@
Lastly, in the adiabatic scheme the final Helium potential is purely repulsive and, therefore, does not allow for bound states. Consequently, the electron spectrum gets contributions only from the continuous final states.
@@ VI The fate of Helium
Fig. 6, panel (a) reports the dynamics of the system for the first picosecond, starting from the equilibrium configuration before the decay with null velocities (i.e., roughly zero temperature). We have studied the dynamics of the system for two different configurations: electronically isolated from the outside, and grounded. The red curves in Fig. 6 correspond to the first instance, in which the system has a total charge of +1, while the black curves correspond to the second instance, where the system is neutralized by gaining an electron from the outside. We find that the Helium detaches in few fs with an energy gain for the system >2>2 eV, as expected (numerical values of these energies for different cases, are reported in the SI, Fig.  SI.1). The oscillations which are visible in Fig. 6 are due to the vibrational excitations induced by the recoil of the Carbon site on which the decay has happened, shown in the structure reported as inset. When the system is neutral (black lines), the behavior is similar, but the energy gain is nearly double, as a consequence of the larger stability of neutral system.
Just after release, the total kinetic energy gets a contribution due to the velocity acquired by the Helium, which is $v_{\rm He}\sim 2.9\times 10^{-5}\simeq 0.088$ Å/fs, for the overall charged case, and $v_{\rm He}\sim 5\times 10^{-5}\simeq 0.15$ Å/fs, when the system is electrically neutral. These values correspond to a Helium kinetic energy of about $1.2$ eV and $3.5$ eV, respectively. As the system evolves, Helium bumps on the graphene layers, and gradually distributes its energy to the substrate over a time scale of ps. This is shown in the upper part of Fig. 6, panel (b), where we also show a prototypical trajectory in the inset. As a consequence, the average energy of the system increases of a few eV.
@@ VII Limitations of our approach and theory systematics
As already explained, to successfully constrain the possible values of the neutrino mass it is of utmost importance to provide a reliable theoretical prediction for the electron’s $\beta$-decay spectrum. The present work is a first step in this direction, as we connect the properties of Tritiated graphene to the shape exhibited by the spectrum. The framework we are currently working in, however, has some limitations, which we now discuss.
@@ VIII Conclusions and Perspectives
The issues raised in the previous section could all be summarized in the consideration that the established and more commonly used theoretical scheme generally rely on some sort of separation of energy or time scales, which in turn allows to decouple a subset of degrees of freedom from others, facilitating the approach to the problem. In this work, conversely, we aimed at addressing a question which is at the connection point of high and low energy scales, overlapping on different time scales, bridging solid state with nuclear physics. Therefore approaches relying on the separation of degrees of freedom, although they can be used as the starting point, needs to be reconsidered and possibly revised.
@@
However, a stronger assumption we made is about the decoupling of the electronic coordinates from the nuclear ones. We used it in two rather opposite declinations: the adiabatic approximation, used before and long after the decay, assuming total adiabaticity of the electrons with respect to the nuclear motion, and the sudden approximation, assuming on the contrary frozen electronic structure. Between these two extremes, a number of theoretical schemes are currently considered to address the relaxation of electronic dynamics during nuclear motion, especially when the latter involves light elements, where quantum effects are more evident. In addition to the evolution of electronic excited states (e.g., within the Time Dependent DFT schemes Marques et al. (2006)) one must also consider the coupled nuclear motion, either within a semi-classical scheme Li et al. (2005), or within the path-integral scheme Kim et al. (1997), if a full quantum accuracy is required. While the latter is still too computationally heavy to address the system of interest, an interesting approach has recently emerged for a quantum perturbative anharmonic treatment with a limited computational cost Monacelli and Mauri (2021).
We also point out that, in order to further reduce the theoretical uncertainty on the correct form of the interaction between Helium and graphene, it would be of great relevance to know, experimentally, with which probability the final state corresponds to He, $\rm He$$^{+}$ or $\rm He$$^{++}$. Such a study for the decay of Tritium on graphene is currently unavailable, although a similar measurement has been performed for the case of molecular Tritium Lin et al. (2020).
The analysis of the model system also requires some consideration. Here we considered only a few possible loading levels, and with symmetric distributions. In a real system, the situation is rather different. While the loading can be somehow controlled and measured, the local Tritium distribution at given loading cannot. On the other hand, our results (and others’ too Delfino et al. (2024); Rossi et al. (2015); Camiola et al. (2015)) indicate that the parameters of the potentials are rather dependent on the local level of loading. Therefore, if one wants to achieve a realistic determination of the parameters of the potential (and in turn of the $\beta$-decay spectrum) one should consider the statistical distribution of possible environments of Tritium/Helium on graphene.
@@
More broadly speaking, a satisfactory formulation of the Tritium $\beta$-decay rate in presence of graphene will have to address a number of points that are still left open by the current analysis. Among these, for example, is the inclusion of long range electrostatic corrections, due to the interaction between the outgoing electron and the rest of the system. These are known to be rather relevant for the decay of Tritium in vacuum, but their relevance in the present context is for now hard to quantify. Moreover, any further refinement of the theoretical predictions should be accompanied by a study of the associated theoretical uncertainties.
All the aspects above will eventually make up the ingredients leading to the determination of the $\beta$-electron spectrum, which should be compared to data to constrain the value of the neutrino mass. To this end, it will be crucial to develop the above proposals, and possibly any other that might be relevant, in the most systematic way possible.